OpenStudy (ledah):

Write the equation of the line parallel to the line 4x + y = -1 and passing through the point (5,0).

OpenStudy (anonymous):

Write the equation in slope intercept form (y = mx + b) where m is the slope. Parallel lines have same slope. So, the new equation will be y = mx + d. Since it passes through point 5,0, you can find the value of d.

OpenStudy (anonymous):

4x + y = -1 => y = -4x -1. So, slope = -4. New line equation is: y = -4x + d Substitute 5,0 for x,y. So, you get: 0 = -4*5 + d => d = 20. So, new line equation is: y = -4x + 20